OpenStudy (anonymous):

f(x) = 1/x^2 Identify any horizontal and vertical asymptotes and identify any holes in the graph.

OpenStudy (anonymous):

What value of x would lead to a zero in the denominator?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

Yes. So if x cannot equal zero, x = 0 must either be a hole or an asymptote. Do you know which it is?

OpenStudy (anonymous):

I think it's an asymptote.

OpenStudy (anonymous):

Yes, exactly. If it were a hole, it would've been canceled out. You've got your answer.

OpenStudy (anonymous):

What about horizontal asymptotes? And how can I tell if there is a hole?

OpenStudy (anonymous):

@greenglasses How can I tell if there's a horizontal asymptote?

OpenStudy (anonymous):

Oh right, I missed the horizontal asymptote part.

OpenStudy (anonymous):

Right, firstly, if there was a hole, you would've cancelled out a value: for example, in the equation: \[f(x)=\frac{ (x+1) }{ (x+1)(x+2)}\]

OpenStudy (anonymous):

x cannot equal -1 or -2, but since you can cancel out x = 1 from the numerator and denominator, x= -1 would be the hole and x = -2 would be the asymptote

OpenStudy (anonymous):

I usually use limits to find a horizontal asymptote- do you know what that is?

OpenStudy (anonymous):

No, but my teacher gave me this rule: Big Top, no asymptote Big Bottom, y=0 BS (I forgot which one it stood for) ex. 4x/3x, y=4/3

OpenStudy (anonymous):

Yes, in this case, the graph has a 'big bottom', and y = 0.

OpenStudy (anonymous):

That's what I thought. May I ask how I would use limits in this situation?

OpenStudy (anonymous):

If you don't know what they are, you probably shouldn't worry about it. Your teacher will teach you eventually if you're in Pre-Calc.

OpenStudy (anonymous):

OK, thank you!